OpenStudy (anonymous):

can someone please explain this to me: A local FM radio station broadcasts at an energy of 6.18E-29 kJ/photon. Calculate the frequency at which it is broadcasting. Frequency = MHz

OpenStudy (aaronq):

Use plank's equation: \(E_{photon}=h\nu\) This will give you the frequency in Hz, so you'll have to convert. Also convert kJ to J before plugging it in.

OpenStudy (anonymous):

okay so i converted and i got 6.18E-26J/photon..... then h=6.626E-34Jsec...so what would i put for E and for v

OpenStudy (aaronq):

E is energy, J is unit of energy. \(\nu\) is frequency, what you're solving for.

OpenStudy (anonymous):

so how would i plug in the value of numbers i got ? sorry im really confused

OpenStudy (aaronq):

yep E=6.18E-26J/photon, h=6.626E-34Jsec. so, \(\nu=\dfrac{6.18E-26J/photon}{6.626E-34Jsec}\)

OpenStudy (anonymous):

v=93268940.54...then divide that by 10^6

OpenStudy (aaronq):

indeed

OpenStudy (anonymous):

how does the photon cancel out when i divde by 10^6 sec-

OpenStudy (aaronq):

i'm not sure which constant cancels it out, i've never written it like that, though it's defined as per photon.

OpenStudy (anonymous):

thank you... could i ask you another question in a couple of mins?

OpenStudy (aaronq):

no problem! sure, post it in a new thread if you don't mind, in case i have to leave.

OpenStudy (anonymous):

will do